# Accelerating Low-field MRI: Compressed Sensing and AI for fast noise-robust imaging

### v1

@@ Discussion
To generate large training, validation, and testing datasets of high-SNR images, data from fastMRI were preprocessed to resemble single-channel $k$-space data acquired on our 6.5 mT MRI scanner  (?). Specifically, multi-channel, fully-sampled raw $k$-space data were sourced from the FastMRI brain training and validation datasets, yielding thousands of 1.5 T / 3 T scans. Coil sensitivity maps were calculated using the ESPIRiT algorithm and SigPy  (?, ?) and used for combining the multi-channel IFFT-reconstructed images into complex-valued, single-channel imagees. Reconstructed brain volumes were converted to 3D $k$-space via FFT and downsampled to the ultra-low field resolution of $64\times 75\times 25$ via $k$-space cropping. Ground-truth reference images were obtained via IFFT reconstruction of those downsampled $k$-space data. Finally, the brain volumes were split into training, validation and test datasets containing 3300, 1100 and 100 volumes respectively.
@@
The training data were further supplemented by 50,000 brain image slices sourced from the Human Connectome Project (HCP)  (?). These brain images were derived from DICOM magnitude images and had synthetic phase maps applied to synthesize complex data acquisition. Individual image slices were downsampled to $75\times 25$ via $k$-space cropping to match the resolution of ultra-low field acquisitions. The synthesis of this HCP dataset is fully described in Reference ?. We emphasize that this dataset was used only for training, i.e. it served as an augmentation method, and not for testing the algorithms.
@@
To analyse the impact of SNR on image-reconstruction performance, for all datasets the downsampled $k$-space data were degraded through the addition of complex-valued additive white Gaussian noise at varying noise power. SNR scales approximately as $B_{0}^{\frac{3}{2}}$  (?). We tested a wide range of SNR values (40 dB range shown in Fig. 2 and Fig. 3) as the variation in SNR expected between clinical and low-field MRI scanners can be very large. Additionally, all images and $k$-space data were linearly scaled such that the 95th percentile of pixel intensity in IFFT reconstructed images was set to 1. This intensity scaling is a standard approach in AI methods for MRI reconstruction.
Ultra-low Field MRI Experiments
@@
Imaging experiments were performed on a 6.5 mT MRI scanner that consists of a biplanar electromagnet-based scanner with a peak gradient strength of 1 mT/m  (?). Image data were acquired in compliance with Institutional Review Board (IRB) approvals with a custom quadrature head coil that has been described previously  (?). Data from this quadrature coil were digitized as two-channel data and then linearly summed after applying a $\pi/2$ phase shift to $k$-space data from the second channel.
A 3D Cartesian balanced-steady state free precession (bSSFP) sequence with TR/TE = 22/11 ms and matrix size $64\times 75\times 25$ (Readout $\times$ Phase Encode 1 $\times$ Phase Encode 2) was programmed with a Tecmag Restone spectrometer and used for data acquisition. These parameters gave a voxel size on the 6.5 mT scanner of 2.5 mm×$\times$3.5 mm×$\times$8 mm (RO×$\times$×PE1$\times$×PE2). The number of excitations (NEX) was varied as described in the Results section.
@@
For prospectively undersampled acquisitions, Poisson disc masks with acceleration factors of $R=2$ and $R=4$ were exported from SigPy and used to populate the phase encoding tables on the spectrometer prior to acquisition. Fully-sampled acquisitions were captured in the same imaging session for comparison of reconstruction methods.
MRI at 3T
@@
High-field imaging data were collected in compliance with IRB approvals from a healthy volunteer on a 3 T Siemens Skyra scanner. A single-channel head coil and a 3D bSSFP sequence (known as True Fast Imaging with Steady-State Free Precession or TRUFI on Siemens scanners) were used for data acquisition. The 3 T bSSFP data shown in Figure 6 were acquired with TR/TE = 4.8 ms/2.4 ms and a matrix size of $256\times 256\times 128$ (Readout $\times$ Phase Encode 1 $\times$ Phase Encode 2) in the anterior-posterior, left-right and superior-inferior directions, respectively. Fully-sampled data were retrospectively downsampled via $k$-space cropping to the same resolution used at ultra-low field ($64\times 75\times 25$).
Fully-sampled data were additionally acquired with gradient-recalled echo (GRE) and Magnetization Prepared Rapid Gradient Echo (MPRAGE) sequences using TR/TE values of 8 ms/2.3 ms and 2300 ms/4.4 ms, respectively. Downsampling was performed using the same $k$-space cropping methodology and the results of reconstruction tests performed using these data are shown in Supplementary Figure 3.
Image Reconstruction
Retrospective undersampling
@@
For retrospective undersampling experiments, 2D masks were created in Python using Sigpy  (?) and then applied to fully-sampled data in both phase encode directions. Poisson-disc masks were created using the corresponding tool in Sigpy, with a $10\times 10$ fully-sampled region added to the centre. Undersampling factors were increased to ensure the reported acceleration factors (2 or 4) were accurate afer addition of this fully-sampled region.
IFFT and CS reconstructions
@@
Fully-sampled data were reconstructed using a conventional IFFT operation. For undersampled data, missing $k$-space data entries were first zero-filled, and images were then reconstructed via IFFT. The zero-filled data were also used as a starting point for $l1$-wavelet CS reconstruction, using the Sigpy package  (?). The CS regularization parameter ($\lambda$) was optimized for each dataset and acceleration factor sperately by performing a grid search over the range of [10-6-10-1]. The iteration number used for the reconstructions was then optimized over the range [1-200].
@@
For experiments using the fastMRI dataset, the optimal $\lambda$ and iteration numbers for CS reconstruction were calculated as the values that minimized the mean NRMSE across 20 brain volumes drawn from the validation set as compared to the ground truth image. Values were reoptimized for each variation in input SNR. For retrospective and prospective experiments on our MRI scanners, the $\lambda$ and iteration values were chosen to minimize NRMSE relative to the fully-sampled image. This parameter re-optimization means that the CS algorithm is being evaluated in terms of its ‘best possible performance’ in each experiment.
AI methods implementation
@@
AUTOMAP: The AUTOMAP reconstruction framework was implemented using supervised learning in Keras and TensorFlow  (?, ?). The AUTOMAP network architecture comprises six trainable layers. The first two layers are fully-connected (dense) layers with hyperbolic tangent activations, which process the flattened input data through hidden layers. The data are then reshaped into a matrix format before passing through three convolutional layers. The first convolutional layer has 128 filters with a 5 × 5 kernel and uses a hyperbolic tangent activation function. The second and third convolutional layers each have 128 filters with 5 × 5 kernels and utilize rectified linear unit (ReLU) activation functions. Finally, a transposed convolutional layer with a single filter and a 5 × 5 kernel generates the final image. This architecture enables the network to learn the spatial decoding process between $k$-space and image space without any prior knowledge. The AUTOMAP network contains 21,925,360 trainable parameters for the $75\times 25$ input size. Visual descriptions of the AUTOMAP architecture were provided in References ?, ?, ?.
@@
Unrolled AI: Our unrolled AI approach to image reconstruction was implemented in PyTorch. The Unrolled architecture consists of 10 iterative soft shrinkage-thresholding blocks. Each block begins with a data fidelity calculation, followed by nonlinear transforms and a soft thresholding operation. The nonlinear transforms include forward and backward operations, each designed as a combination of two convolutional layers split by a rectified linear unit (ReLU). A skip connection is incorporated to form a residual block, aiding in network training. A data consistency operation is included at the end of each block (except for the final block) to ensure $k$-space consistency with the measured $k$-space samples. This data consistency layer replaces AI-generated $k$-space data with the original data at sampled points. In the last iteration, the data consistency block is omitted to allow the convolutional layers to denoise the original low SNR $k$-space data that was reinjected in previous iterations. As implemented, the Unrolled network has 98,760 parameters, a value that is not dependent on the input size. The full Unrolled network architecture is detailed in Reference ?.
@@
Training: Both image reconstruction networks were trained using the same training corpus of fastMRI and HCP data described above. To minimize graphical processing unit (GPU) memory requirements, networks were trained to reconstruct 2D, rather than 3D, data  (?). Hence, the training corpus was utilized as a stack of independent 2D $75\times 25$ brain image/masked $k$-space pairs. Data were further augmented by adding random amounts of synthetic Gaussian noise to the input during training, with a variable magnitude range that mimicked the SNR expected during implementation. Individual networks were trained for each undersampling mask.
A custom loss function was defined to calculate the mean square error (MSE) with the background region outside the brain/head masked. This masking encourages the neural networks to focus on improving image structure within the brain rather than denoising the background  (?). Masking was performed in regions where the normalized reference images had a signal magnitude less than 0.005.
Each network was trained for up to 100 epochs. AUTOMAP training utilized a learning rate of 0.0001 and the RMSprop optimizer. Unrolled AI networks were trained with a learning rate of 0.001 and the Adam optimizer.
Two separate AUTOMAP networks were trained for each reconstruction task, one to reconstruct real image data and one to reconstruct imaginary image data. AUTOMAP networks received the same complex $k$-space input data but were shown different (real or imaginary) reference images.
@@
Loss functions. Additional experiments were performed to test the effects of perceptual loss functions during training  (?). Results from these experiments are reported in Supplementary Figure 2. To train the unrolled network, we utilized the Learned Perceptual Image Patch Similarity (LPIPS) loss available for PyTorch at https://github.com/richzhang/PerceptualSimilarity. To train AUTOMAP, we ported this LPIPS loss to TensorFlow. Our implementation is available at:
https://github.com/Image-X-Institute/lpips_torch2tf  (?).
@@
Inference. All reconstructions presented in the Results section were performed on 3D $k$-space data. To address GPU memory limitations, a 1D FFT was first applied along the fully-sampled readout dimension of $k$-space to yield a stack of hybrid $k$-space data. The trained networks were then applied to the hybrid $k$-space data to perform reconstruction in undersampled phase encode directions via inference. In the case of AUTOMAP, inference was performed separately with real and imaginary networks before outputs were combined into a complex image.
Metric Calculations

### v2

@@ 2.1.1 Data preprocessing for in silico experiments
To generate large training, validation, and testing datasets of high-SNR images, brain data from the fastMRI database were preprocessed to resemble single-channel $k$-space data acquired on our 6.5 mT MRI scanner [1, 43]. This raw data was made publicly available by the NYU fastMRI Initiative specifically for the purpose of exploring whether machine learning can aid in the reconstruction of medical images [43, 44]. In this experiment, multi-channel, fully-sampled raw $k$-space data were sourced from the fastMRI brain training and validation datasets, yielding thousands of 1.5 T / 3 T scans. Coil sensitivity maps were calculated using the ESPIRiT algorithm and SigPy [45, 46] and used for combining the multi-channel inverse fast Fourier transform (IFFT) reconstructed images into complex-valued, single-channel images. Reconstructed brain volumes were converted to 3D $k$-space via fast Fourier transform (FFT) and downsampled to the ultra-low field resolution of $64\times 75\times 25$ via $k$-space cropping. Ground-truth reference images were obtained via IFFT reconstruction of these downsampled $k$-space data. Finally, the brain volumes were split into training, validation and test datasets containing 3300, 1100 and 100 volumes respectively.
@@
The training dataset was further supplemented with 50,000 brain image slices sourced from the Human Connectome Project (HCP) [47]. These brain images were derived from DICOM magnitude images and had synthetic phase maps applied to synthesize complex data acquisition. Individual image slices were downsampled to $75\times 25$ via $k$-space cropping to match the resolution of ultra-low field acquisitions. The synthesis of this HCP dataset is fully described in Reference 36. We emphasize that this dataset was used only for training, i.e. it served as an augmentation method, and not for testing the algorithms.
@@
To analyse the impact of SNR on image-reconstruction performance, for all datasets the downsampled $k$-space data were degraded through the addition of complex-valued additive white Gaussian noise at varying noise power. The quality of images reconstructed from this degraded $k$-space data was subsequently compared to reference images reconstructed from the original fully-sampled and noise-free data using the methods described below. As SNR scales approximately as $B_{0}^{\frac{3}{2}}$ [48], we tested a wide range of SNR values (40 dB range shown in Fig. 2 and Fig. 3) as the variation in SNR expected between clinical and low-field MRI scanners can be very large. Additionally, images and $k$-space data were linearly scaled such that the 95th percentile of pixel intensity in IFFT reconstructed images was set to 1.
2.2 MRI Data Acquisition
Experimental data in this study were acquired at the ultra-low field of 6.5 mT and the standard clinical field strength of 3 T. All data were acquired with informed consent in compliance with Institutional Review Board (IRB) approvals. MRI data were acquired under Mass General Brigham protocol number 2022P002344, which was approved on July 10, 2024.
2.2.1 Ultra-low Field MRI Experiments
@@
Imaging experiments were performed on a 6.5 mT MRI scanner that consists of a biplanar electromagnet-based scanner with a peak gradient strength of 1 mT/m [1]. Image data were acquired with a custom quadrature head coil that has been described previously [49]. Data from this quadrature coil were digitized as two-channel data and then linearly summed after applying a $\pi/2$ phase shift to $k$-space data from the second channel. The 6.5 mT MRI system includes an RF-shielded enclosure, as described in Reference 1, which reduces environmental electromagnetic interference during acquisition.
[…]
@@
For prospectively undersampled acquisitions, Poisson disc masks with acceleration factors of $R=2$ and $R=4$ were exported from SigPy and used to populate the phase encoding tables on the spectrometer prior to acquisition. Fully-sampled acquisitions were captured in the same imaging session for comparison of reconstruction methods. All scans were acquired independently with $\mathrm{NEX}=48$, so that the accelerated acquisitions were completed in half or a quarter of the fully-sampled time for $R=2$ and $R=4$ masks, respectively.
2.2.2 MRI at 3T
@@
High-field imaging data were collected in compliance with IRB approvals from a healthy volunteer on a 3 T Siemens Skyra scanner. A single-channel head coil and a 3D bSSFP sequence were used for data acquisition. The 3 T bSSFP data shown in Figure 6 were acquired with TR/TE = 4.8 ms/2.4 ms and a matrix size of $256\times 256\times 128$ (Readout $\times$ Phase Encode 1 $\times$ Phase Encode 2) in the anterior-posterior, left-right and superior-inferior directions, respectively. Fully-sampled data were retrospectively downsampled via $k$-space cropping to the same resolution used at ultra-low field ($64\times 75\times 25$). Fully-sampled data were additionally acquired with gradient-recalled echo (GRE) and Magnetization Prepared Rapid Gradient Echo (MPRAGE) sequences using TR/TE values of 8 ms/2.3 ms and 2300 ms/4.4 ms, respectively.
2.3 Image Reconstruction
2.3.1 Retrospective undersampling
@@
For retrospective undersampling experiments, 2D masks were created in Python using Sigpy [46] and then applied to fully-sampled data in both phase encode directions. Poisson-disc masks were created using the ‘Poisson-disc’ tool in Sigpy, with a $10\times 10$ fully-sampled region added to the center. The reported acceleration factors (2 or 4) were calculated after addition of this fully-sampled region. Additional masks used for undersampling are shown with the accompanying results where relevant.
2.3.2 IFFT and CS reconstructions
@@
Fully-sampled data were reconstructed using a conventional IFFT operation. For undersampled data, missing $k$-space data entries were first zero-filled, and images were then reconstructed via IFFT. The 95th percentile intensity value of the zero-filled images was linearly scaled to 1 and the resultant scaling factor applied to the undersampled $k$-space as a starting point for subsequent $l1$-wavelet CS reconstruction using the Sigpy package [46]. The CS regularization parameter ($\lambda$) was optimized for each dataset and acceleration factor separately by performing a grid search over the range of [10-6-10-1]. The iteration number used for the reconstructions was then optimized over the range [1-200].
@@
For experiments using the fastMRI dataset, the optimal $\lambda$ and iteration numbers for CS reconstruction were calculated as the values that minimized the mean normalized root-mean-square error (NRMSE) across 20 brain volumes drawn from the validation set as compared to the ground truth image. Values were re-optimized for each variation in input SNR. For retrospective and prospective experiments on our MRI scanners, the $\lambda$ and iteration values were chosen to minimize NRMSE relative to the fully-sampled image. This parameter re-optimization means that the CS algorithm is being evaluated in terms of its ‘best possible performance’ in each experiment.
2.3.3 DL reconstruction
We analyzed the performance of three DL reconstruction frameworks: (i) a data-driven framework trained on the forward encoding model known as AUTOMAP [50], (ii) a physics-guided unrolled neural network [51, 52] and (iii) a Swin Transformer cascade.
@@ 2.3.3 DL reconstruction
AUTOMAP: The AUTOMAP reconstruction framework was implemented using supervised learning in Keras and TensorFlow [36, 53]. The AUTOMAP network architecture comprises six trainable layers. The first two layers are fully-connected (dense) layers with hyperbolic tangent activations, which process the flattened input data through hidden layers. The data are then reshaped into a matrix format before passing through three convolutional layers. The first convolutional layer has 128 filters with a 5 × 5 kernel and uses a hyperbolic tangent activation function. The second and third convolutional layers each have 128 filters with 5 × 5 kernels and utilize rectified linear unit (ReLU) activation functions. Finally, a transposed convolutional layer with a single filter and a 5 × 5 kernel generates the final image. This architecture enables the network to learn the spatial decoding process between $k$-space and image space without any prior knowledge. The AUTOMAP network contains 21,925,360 trainable parameters for the $75\times 25$ input size. TensorFlow was utilised to ensure that the AUTOMAP architecture was implemented in the same fashion as originally published in Reference 50. Visual descriptions of the AUTOMAP architecture were provided in References 50, 36 and 53.
@@
Unrolled DL: Our unrolled DL approach to image reconstruction was implemented in PyTorch. The Unrolled architecture consists of 10 iterative soft shrinkage-thresholding blocks. Each block begins with a data fidelity calculation, followed by nonlinear transforms and a soft thresholding operation. The nonlinear transforms include forward and backward operations, each designed as a combination of two convolutional layers split by a rectified linear unit (ReLU). A skip connection is incorporated to form a residual block, aiding in network training. A data consistency (DC) operation is included at the end of each block (except for the final block) to ensure $k$-space consistency with the measured $k$-space samples. This DC layer replaces DL-generated $k$-space data with the original data at sampled points. In the last iteration, the DC block is omitted to allow the convolutional layers to de-noise the original low SNR $k$-space data that was re-injected in previous iterations. As implemented, the Unrolled network has 98,760 parameters, a value that is not dependent on the input size. The full Unrolled network architecture is detailed in Reference 52.
Swin Cascade: The Swin Transformer cascade was implemented in PyTorch using transformer blocks adapted from the SwinIR (Swin Transformer for Image Restoration) framework [54, 55]. The model comprises six cascaded stages, each operating in the image domain and based on residual Swin Transformer layers with shifted window self-attention. Each stage refines the reconstruction using these SwinIR-based transformer blocks. The output is passed through an FFT to update $k$-space before DC. The DC operation was implemented by replacing predicted $k$-space values at sampled locations with the original acquired measurements. This architecture enables the network to process data efficiently with physics constraints, while enabling full denoising in the last stage. The total number of trainable parameters was 8,335,032 for the training input size used. Although this structure mirrors the image–data-consistency staging of our convolutional unrolled DL network, we follow standard terminology and keep the transformer variant labelled a Swin Cascade—similar to other work that explored transformers for high-field MRI [55, 56]—while reserving the term unrolled for architectures that explicitly unfold a specific iterative optimizer using convolutional layers. Additional details of the Swin Transformer design and its application in a Swin Cascade are provided in References 54 and 55.
@@
Training: All image reconstruction networks were trained using the same training corpus of fastMRI and HCP data described above, which consisted of more than 250,000 image slices. To minimize graphical processing unit (GPU) memory requirements, networks were trained to reconstruct 2D, rather than 3D, data [57]. Hence, the training corpus was utilized as a stack of independent 2D $75\times 25$ brain image/masked $k$-space pairs. Data were further augmented by adding random amounts of synthetic Gaussian noise to the input during training, with a variable magnitude range that mimicked the SNR expected during implementation. Individual networks were trained for each undersampling mask. Hyperparameters such as learning rate, and number of unrolls or cascades were selected via manual tuning on a validation subset, balancing training stability, reconstruction fidelity, and GPU memory constraints.
A custom loss function was defined to calculate the mean square error (MSE) after masking the background region outside the brain/head. This masking encourages the neural networks to focus on improving image structure within the brain rather than de-noising the background [58]. Masking was performed in regions where the normalized reference images had a signal magnitude less than 0.005. All networks were trained for up to 100 epochs or 100 hours, whichever came first. While the Unrolled and AUTOMAP models were trained for the full 100 epochs, the Swin Cascade models were trained for 12 epochs, which was sufficient for convergence in all cases. Although the Swin Cascade has ~8.3 M learnable parameters—fewer than the ~20 M in AUTOMAP—it stacks many more layers and performs multi-head windowed self-attention in every block. As a result, the Swin model uses substantially more GPU memory and takes longer to train than AUTOMAP, despite its smaller parameter count [59, 60]. AUTOMAP training utilized a learning rate of 0.0001 and the RMSprop optimizer. Unrolled DL networks were trained with a learning rate of 0.001 and the Adam optimizer. Swin Cascade networks were trained with a learning rate of 0.0001 and the Adam optimizer. Two separate AUTOMAP networks were trained for each reconstruction task, one to reconstruct real image data and one to reconstruct imaginary image data. AUTOMAP networks received the same complex $k$-space input data but were shown different (real or imaginary) reference images.
@@
Additional experiments were performed to test the effects of perceptual loss functions during training [61]. To train the unrolled network, we utilized the Learned Perceptual Image Patch Similarity (LPIPS) loss available for PyTorch at https://github.com/richzhang/PerceptualSimilarity. To train AUTOMAP, we ported this LPIPS loss to TensorFlow. Our implementation is available at:
https://github.com/Image-X-Institute/lpips_torch2tf [62].
@@
Inference. All reconstructions of test data were performed on 3D $k$-space data. To address GPU memory limitations, a 1D FFT was first applied along the fully-sampled readout dimension of $k$-space to yield a stack of hybrid $k$-space data. The trained networks were then applied to the hybrid $k$-space data to perform reconstruction in undersampled phase encode directions via inference. In the case of AUTOMAP, inference was performed separately with real and imaginary networks before outputs were combined into a complex image.
2.3.4 Metric Calculations and Hidden Noise